# Beyond the Electric Dipole Approximation: Electric and Magnetic Multipole Contributions Reveal Biaxial Water Structure from SFG Spectra at the Air-Water Interface

### v2

@@ II.4 Spatially Resolved Second-Order Response Profile of the Stretch Band
Figure 4: Second-order response profile $\tilde{s}^{(2)}_{ijk}\left(z\right)$ of the OH-stretch band as defined in Equation (7) and the decomposition into its multipole components $\tilde{s}^{(2,\beta)}_{ijk}\left(z\right)$ as defined in Equation (17). A snapshot of the simulation box is shown in (a). The mass-density profiles with respect to $z_{\mathrm{GDS}}$ and $z_{\mathrm{WCS}}$ are presented in (b). Slices of $\tilde{s}^{(2)}_{ijk}\left(z\right)$ are shown for selected positions in (c) & (d) relative to $z_{\mathrm{GDS}}$ and in (e) & (f) relative to $z_{\mathrm{WCS}}$, colored vertical lines in (a), (b), (g)-(j) indicate the positions. The profiles $\tilde{s}^{(2)}_{ijk}\left(z\right)$ are shown relative to $z_{\mathrm{WCS}}$ in (g) & (h) and relative to $z_{\mathrm{GDS}}$ in (i) & (j). The profile relative to $z_{\mathrm{GDS}}$ is dissected into its molecular multipole components $\tilde{s}^{(2,\mathrm{D})}_{ijk}\left(z\right)$ and $\tilde{s}^{(2,\mathrm{Q})}_{ijk}\left(z\right)$ in (k) & (l) and (m) & (n), respectively. The spectra are red-shifted by $166\text{\,}\mathrm{c}\mathrm{m}^{-1}$.
We present the imaginary part of the second-order response profile $\tilde{s}^{(2)}_{ijk}(z)$ in the OH-stretch frequency region, defined in Equation (7), together with a decomposition into the electric dipole and electric quadrupole contribution according to Equation (17) in Figure 4. We map the waters center of mass positions in the calculation of $\tilde{s}^{(2)}_{ijk}(z)$ relative to the non-planar Willard-Chandler surface $z_{\mathrm{WCS}}$ [69] in Figure 4 (e)-(h) and relative to the planar Gibbs dividing surface $z_{\mathrm{GDS}}$ in Figure 4 (c) & (d) and (i)-(n). We show a snapshot of the simulation box in Figure 4 (a). In Figure 4 (b), we present the mass-density profiles relative to $z_{\mathrm{GDS}}$ and $z_{\mathrm{WCS}}$. In the latter presentation the density profile strongly oscillates, while in the laboratory frame these oscillations are washed out due to the intrinsic roughness of the air-water interface [70]. Slices of $\tilde{s}^{(2)}_{ijk}(z)$ at selected positions $z-z_{\mathrm{GDS}}$ and $z-z_{\mathrm{WCS}}$ (denoted by the colored lines in (b)) are presented in Figure 4 (c) & (d) and (e) & (f). Close to the vapor region at $z-z_{\mathrm{GDS/WCS}}=$0.5\text{\,}\mathrm{\SIUnitSymbolAngstrom}$\$ (orange lines), almost only the free OH stretch peak is visible at $3700\text{\,}\mathrm{c}\mathrm{m}^{-1}$. At $z-z_{\mathrm{GDS/WCS}}=$-0.5\text{\,}\mathrm{\SIUnitSymbolAngstrom}$\$ (purple lines), $\tilde{s}^{(2)}_{ijk}(z)$ qualitatively agrees with the integrated spectra presented in Figure 1 (g) & (h). Finally, the free OH stretch contributions vanish near the bulk region at $z-z_{\mathrm{GDS/WCS}}=$-1.5\text{\,}\mathrm{\SIUnitSymbolAngstrom}$\$ (turquoise lines), and we only see the shoulder at $3600\text{\,}\mathrm{c}\mathrm{m}^{-1}$ and the negative component at $3500\text{\,}\mathrm{c}\mathrm{m}^{-1}$. As shown in Figure 4 (h), $\tilde{s}^{(2)}_{zzz}(z)$ relative to $z_{\mathrm{WCS}}$ almost perfectly follows the oscillations of the mass density profile in Figure 4 (b). The second-order response profile in the laboratory frame $\tilde{s}^{(2)}_{ijk}(z)$ is presented in Figure 4 (c) & (d) and (i) & (j). Here, both the free OH stretch band at $3700\text{\,}\mathrm{c}\mathrm{m}^{-1}$ and the negative contribution at $3500\text{\,}\mathrm{c}\mathrm{m}^{-1}$ are located at approximately $z\approx z_{\mathrm{GDS}}$. However, the shoulder is created below $z_{\mathrm{GDS}}$, and the free-OH contributions reaches slightly more into the vapor region. The profiles of the second-order response of the electric dipole contribution $\tilde{s}^{(2,\mathrm{D})}_{ijk}(z)$ are presented in Figure 4 (k) & (l). These qualitatively agree with previously published depth-resolved SFG spectra [71, 72, 73, 74, 21, 75, 76]. The second-order electric dipole contributions in the stretching band (Fig. 4 (k) & (l)) and in the bending band (Fig. 2 (f)) undergo significant frequency shifts as a function of position $z$ in the interfacial layer. As shown elsewhere [77], when going from vapor to liquid, frequency shifts arise from competing effects of non-Markovian friction (which causes blue-shifting) and potential broadening (which causes red-shifting). The former dominates in the bending band. The electric quadrupole profile $\tilde{s}^{(2,\mathrm{Q})}_{ijk}(z)$ is presented in Figure 4 (m) & (n) and is dominated by a broad positive peak around $z\approx z_{\mathrm{GDS}}$. Additionally, we observe a negative contribution closer to the bulk region and a positive contribution closer to the vapor region at the frequency of the free OH vibrations. It is important to note that the electric quadrupole profile $\tilde{s}^{(2,\mathrm{Q})}_{ijk}(z)$ strongly depends on position. Consequently, the electric quadrupole contribution to the SFG spectra, $\tilde{S}^{(2,\mathrm{Q})}_{ijk}$ as defined in Equation (18), does report on interfacial water structure for a non-vanishing wave vector mismatch $\Delta k_{z}$, which is an interesting venue for future experimental investigation.
@@ II.5 Spatially Resolved Linear Dielectric and Absorption Profile
Figure 5: Linear dielectric and absorption profiles at the air water interface. (a): Snapshot of the simulation. (b): The density profile is compared to the real part of the optical dielectric profile. (c) & (d): The linear absorption profiles $\omega\tilde{\varepsilon}^{\prime\prime}_{xx}(z)$ and $\omega\tilde{\varepsilon}^{\prime\prime}_{zz}(z)$ in the IR frequency range are shown as a 2D plot. (e)-(j): The imaginary part of the linear absorption profile is shown at the same positions relative to the Gibbs dividing surface as in Figure 4. The xxxx-component defined in Equation (25), the zzzz-component defined in Equation (24), and the isotropic component defined in (26) are presented in (e), (f), and (g), respectively. Additionally, the predicted linear absorption spectrum of bulk water and the experimental spectrum [78] are shown in (g). (h)-(j): Results in the stretch region. The bending band in (e)-(g) is redshifted by $28\text{\,}\mathrm{c}\mathrm{m}^{-1}$ and the stretching band in (h)-(j) is red-shifted by $166\text{\,}\mathrm{c}\mathrm{m}^{-1}$.
As mentioned before, the $z$-independent external field amplitude $\mathcal{F}^{\alpha}_{z}$ corresponds to the D-field, while $\mathcal{F}^{\alpha}_{x/y}$ corresponds to the E-field. The relationship between D- and E-fields is determined by the dielectric profile [79, 80, 81, 82, 83, 84, 85]. The real part of the dielectric profile $\tilde{\varepsilon}_{xx}^{\mathrm{VIS}}(z)$ (red) and $\tilde{\varepsilon}_{zz}^{\mathrm{VIS}}(z)$ (blue) at optical frequencies is presented in Figure 5 (b) and has an almost identical shape to the mass density profile (black). Thus, in contrast to the static or IR case [80, 86], the anisotropy of the tensorial interfacial response for optical frequencies is rather small, as demonstrated in the inset. The bulk plateau value of $1.77$ agrees well with the experimental value of $1.78$ at room temperature under atmospheric pressure [87]. We present the imaginary part of the position and frequency-dependent dielectric profiles, $\tilde{\varepsilon}^{\mathrm{IR}}_{xx}(z)$ and $\tilde{\varepsilon}^{\mathrm{IR}}_{zz}(z)$ extracted from molecular dynamics simulations using Equations (24) and (25) in Figure 5 (c) & (d). We observe that the dielectric profile $\tilde{\varepsilon}^{\mathrm{IR}}_{ij}(z)$ varies drastically in the interface region and is significantly anisotropic. We see that the change in the parallel component of the dielectric profile $\tilde{\varepsilon}^{\mathrm{IR}}_{xx}\left(z\right)$ (Fig. 5 (c) & (e) & (h)) is mainly characterized by a frequency-independent decrease of intensity across the interface, without significant changes in the line shape. The transition of the perpendicular component $\tilde{\varepsilon}_{zz}(z)$ (Fig. 5 (d) & (f) & (i)) is more complex. The OH-stretch band, which is in bulk centered around $3500\text{\,}\mathrm{c}\mathrm{m}^{-1}$, undergoes a blue shift from the bulk to vapor until only the free OH-stretch part at $3700\text{\,}\mathrm{c}\mathrm{m}^{-1}$ survives. The libration band [88, 89], between $250\text{\,}\mathrm{c}\mathrm{m}^{-1}$ and $1000\text{\,}\mathrm{c}\mathrm{m}^{-1}$, is more pronounced in $\tilde{\varepsilon}^{\mathrm{IR}}_{zz}(z)$ than in $\tilde{\varepsilon}^{\mathrm{IR}}_{xx}(z)$, as can be seen in Figures 5 (e) & (f). This can be explained by the anisotropic orientation of the molecular dipole moments, and likely results from a higher concentration of planar-oriented molecules near the interface (Fig. 3 (e)). We observe that the interfacial and the bulk spectra are indistinguishable at $z-z_{\mathrm{GDS}}=$-5.5\text{\,}\mathrm{\SIUnitSymbolAngstrom}$\$ in Figure 5 (g) & (j). Hence, the second-order response in Figures 2 & 4, as well as the linear response in Figure 5, and the orientational anisotropy in Figure 3, are bulk-like at this depth. We compare with experiments [78] (black) in Figure 5 (g) & (j), where we observe that the differences between simulation and experiments are rather small in the wavenumber range presented in Figure 5 (g). Especially, the difference between the experimental bending peak and simulation is small. In the OH-stretch region in Figure 5 (j), we observe good agreement at high-frequencies ($3500\text{\,}\mathrm{c}\mathrm{m}^{-1}$\$) and an underestimation of the experimental IR intensity at lower frequencies.
@@ III Conclusion
Accounting for higher-order multipole contributions is essential for interpreting SFG spectra and leads to quantitative agreement between simulations and experimental data. In the bending band, the electric dipole contribution is almost negligible due to the small frequency shift between mutually canceling positive and negative SFG signals from water at different depths, and has an amplitude comparable to the previously neglected magnetic dipole contribution. As a result, the total SFG signal is dominated by the electric quadrupole contribution. The shoulder appearing at $3600\text{\,}\mathrm{c}\mathrm{m}^{-1}$ in the stretch band has been attributed to orientational anisotropy [90, 91], or to a combination mode involving intermolecular coupling [92], as summarized in the review by Tang et al. [93]. Our results demonstrate that the shoulder primarily originates from electric quadrupole contributions, which do not reflect interfacial structure but rather depend on bulk properties for vanishing wave vector mismatch $\Delta k_{z}=0$. After subtracting higher-order multipoles, the anisotropic orientation of interfacial molecules determines the SFG signal which reveals strong biaxial water ordering at the air-water interface. This biaxial structure is composed of a layering of essentially three orientational species, namely an inwards-oriented, a planar-oriented, and an outwards-oriented species.
In conclustion, since SFG spectra are significantly influenced by bulk-specific multipole effects, it is crucial to account for and subtract these contributions to interpret experimental spectra in terms of interfacial orientations. This can be achieved by subtracting a reference spectrum with the same bulk medium [50], or by using theoretical predictions, as those provided in this work.

### v3

@@ II.4 Spatially Resolved Second-Order Response Profile of the Stretch Band
Figure 4: Second-order response profile $\tilde{s}^{(2)\prime\prime}_{ijk}\left(z\right)$ of the OH-stretch band as defined in Equation (8) and the decomposition into its multipole components $\tilde{s}^{(2,\beta)\prime\prime}_{ijk}\left(z\right)$ as defined in Equation (20). A snapshot of the simulation box is shown in (a). The mass-density profiles with respect to $z_{\mathrm{GDS}}$ and $z_{\mathrm{WCS}}$ are presented in (b). Slices of $\tilde{s}^{(2)\prime\prime}_{ijk}\left(z\right)$ are shown for selected positions in (c) & (d) relative to $z_{\mathrm{GDS}}$ and in (e) & (f) relative to $z_{\mathrm{WCS}}$, colored vertical lines in (a), (b), (g)-(j) indicate the positions. The profiles $\tilde{s}^{(2)\prime\prime}_{ijk}\left(z\right)$ are shown relative to $z_{\mathrm{WCS}}$ in (g) & (h) and relative to $z_{\mathrm{GDS}}$ in (i) & (j). The profile relative to $z_{\mathrm{GDS}}$ is dissected into its molecular multipole components $\tilde{s}^{(2,\mathrm{D})\prime\prime}_{ijk}\left(z\right)$ and $\tilde{s}^{(2,\mathrm{Q})\prime\prime}_{ijk}\left(z\right)$ in (k) & (l) and (m) & (n), respectively. All spectra are red-shifted by $166\text{\,}\mathrm{c}\mathrm{m}^{-1}$.
We present the imaginary part of the second-order response profile $\tilde{s}^{(2)\prime\prime}_{ijk}(z)$ in the OH-stretch frequency region, defined in Equation (8), together with a decomposition into the electric dipole and electric quadrupole contributions according to Equation (20) in Figure 4. We map the waters center of mass positions in the calculation of $\tilde{s}^{(2)}_{ijk}(z)$ relative to the non-planar Willard-Chandler surface $z_{\mathrm{WCS}}$ [134] in Figure 4 (e)-(h) and relative to the planar Gibbs dividing surface $z_{\mathrm{GDS}}$ in Figure 4 (c) & (d) and (i)-(n). We show a snapshot of the simulation box in Figure 4 (a). In Figure 4 (b), we present the mass-density profiles relative to $z_{\mathrm{GDS}}$ and $z_{\mathrm{WCS}}$. In the latter presentation the density profile strongly oscillates, while in the laboratory frame these oscillations are washed out due to the intrinsic roughness of the air-water interface [109].
Slices of $\tilde{s}^{(2)\prime\prime}_{ijk}(z)$ at selected positions $z-z_{\mathrm{GDS}}$ and $z-z_{\mathrm{WCS}}$ (denoted by the colored lines in (b)) are presented in Figure 4 (c) & (d) and (e) & (f). Close to the vapor at $z-z_{\mathrm{GDS/WCS}}=$0.5\text{\,}\mathrm{\SIUnitSymbolAngstrom}$\$ (orange lines), almost only the free OH stretch peak is visible at $3700\text{\,}\mathrm{c}\mathrm{m}^{-1}$. At $z-z_{\mathrm{GDS/WCS}}=$-0.5\text{\,}\mathrm{\SIUnitSymbolAngstrom}$\$ (purple lines), $\tilde{s}^{(2)\prime\prime}_{ijk}(z)$ resembles the integrated spectra presented in Figure 1 (g) & (h). Finally, the free OH stretch contributions vanish near the bulk at $z-z_{\mathrm{GDS/WCS}}=$-1.5\text{\,}\mathrm{\SIUnitSymbolAngstrom}$\$ (turquoise lines), and we only see the shoulder at $3600\text{\,}\mathrm{c}\mathrm{m}^{-1}$ and the negative component at $3500\text{\,}\mathrm{c}\mathrm{m}^{-1}$. As shown in Figure 4 (h), $\tilde{s}^{(2)\prime\prime}_{zzz}(z)$ relative to $z_{\mathrm{WCS}}$ almost perfectly follows the oscillations of the mass density profile in Figure 4 (b).
The imaginary part of the second-order response profile in the laboratory frame $\tilde{s}^{(2)\prime\prime}_{ijk}(z)$ is presented in Figure 4 (c) & (d) and (i) & (j). Here, both the free OH stretch band at $3700\text{\,}\mathrm{c}\mathrm{m}^{-1}$ and the negative contribution at $3500\text{\,}\mathrm{c}\mathrm{m}^{-1}$ are located at approximately $z\approx z_{\mathrm{GDS}}$. However, the shoulder is created below $z_{\mathrm{GDS}}$, and the free-OH contributions reaches slightly more into the vapor region. The profiles of the second-order response of the electric dipole contribution $\tilde{s}^{(2,\mathrm{D})\prime\prime}_{ijk}(z)$ are presented in Figure 4 (k) & (l). These qualitatively agree with previously published depth-resolved SFG spectra [86, 62, 83, 58, 39, 40, 32].
The second-order electric dipole contributions to the SFG signal in the stretching band (Fig. 4 (k) & (l)) and to the susceptibility in the bending band (Fig. 2 (f)) undergo significant frequency shifts as a function of position $z$ in the interfacial layer. As shown elsewhere [24], when going from vapor to liquid, frequency shifts arise from competing effects of non-Markovian friction (which causes blue-shifting) and potential broadening (which causes red-shifting). The former dominates in the bending band, as seen in Figure 2 (f), while the latter dominates in the stretch band, as seen in Figure 4 (k) & (l).
The electric quadrupole profile $\tilde{s}^{(2,\mathrm{Q})\prime\prime}_{ijk}(z)$ is presented in Figure 4 (m) & (n) and is dominated by a broad positive peak around $z\approx z_{\mathrm{GDS}}$. Additionally, we observe a negative contribution closer to the bulk region and a positive contribution closer to the vapor region at the frequency of the free OH vibrations. It is important to note that the electric quadrupole profile $\tilde{s}^{(2,\mathrm{Q})\prime\prime}_{ijk}(z)$ strongly depends on position. Consequently, the electric quadrupole contribution to the SFG spectra, $\tilde{S}^{(2,\mathrm{Q})}_{ijk}$ as defined in Equation (21), could be used to report on interfacial water structure for a non-vanishing wave vector mismatch $\Delta k_{z}$, which is an interesting venue for future experimental investigation.
While the bending band probes mostly the biaxial interfacial water ordering, as shown in Section II.3, the stretch band is instead more sensitive to the local hydrogen-bonding environment, which varies rather abruptly across the interface [127, 31]. Therefore, the second-order electric-dipole profile of the stretch band is confined to a narrow region around the Gibbs dividing surface, whereas the bending profile extends slightly deeper into the bulk, as seen by comparing Figure 4 (k) with Figure 2 (f).
@@ II.5 Spatially Resolved Linear Dielectric and Absorption Profile
Figure 5: Linear optical dielectric and IR absorption profiles at the air water interface. (a): Snapshot of the simulation. (b): The water mass density profile is compared to the real parts of the tensorial optical dielectric profiles. (c) & (d): The linear absorption profiles $\omega\tilde{\varepsilon}^{\prime\prime}_{xx}(z)$ and $\omega\tilde{\varepsilon}^{\prime\prime}_{zz}(z)$ in the IR and low THz frequency range are shown as a 2D plot. (e)-(j): The imaginary part of the linear absorption profile is shown at the same positions relative to the Gibbs dividing surface as in Figure 4. The xxxx-component defined in Equation (28), the zzzz-component defined in Equation (27), and the isotropic average defined in (29) are presented in (e), (f), and (g), respectively. Additionally, the predicted linear absorption spectrum of bulk water and the experimental spectrum [16] are shown in (g). (h)-(j): Results in the stretch region. The bending band in (e)-(g) is redshifted by $28\text{\,}\mathrm{c}\mathrm{m}^{-1}$ and the stretching band in (h)-(j) is red-shifted by $166\text{\,}\mathrm{c}\mathrm{m}^{-1}$.
As mentioned before, the $z$-independent external field amplitude $\mathcal{F}^{\alpha}_{z}$ corresponds to the D-field, while $\mathcal{F}^{\alpha}_{x/y}$ corresponds to the E-field. The relationship between D- and E-fields is determined by the dielectric profile [120, 19, 20, 21, 47, 75, 76]. The real parts of the dielectric profiles $\tilde{\varepsilon}_{xx}^{\mathrm{VIS}\prime}(z)$ (red) and $\tilde{\varepsilon}_{zz}^{\mathrm{VIS}\prime}(z)$ (blue) at optical frequencies are presented in Figure 5 (b) and have an almost identical shape to the mass density profile (black). Thus, in contrast to the static or THz case [19, 12], the anisotropy of the tensorial interfacial response for optical frequencies is rather small, as demonstrated in the inset. The bulk plateau value of $1.77$ agrees well with the experimental value of $1.78$ at room temperature under atmospheric pressure [55].
We present the imaginary part of the position and frequency-dependent dielectric profiles, $\tilde{\varepsilon}^{\mathrm{IR}\prime\prime}_{xx}(z)$ and $\tilde{\varepsilon}^{\mathrm{IR}\prime\prime}_{zz}(z)$ extracted from molecular dynamics simulations using Equations (27) and (28) in Figure 5 (c) & (d). We observe that the dielectric profile $\tilde{\varepsilon}^{\mathrm{IR}\prime\prime}_{ij}(z)$ varies drastically in the interface region and is significantly anisotropic. We see that the parallel component of the dielectric profile $\tilde{\varepsilon}^{\mathrm{IR}\prime\prime}_{xx}\left(z\right)$ (Fig. 5 (c) & (e) & (h)) is characterized by an almost frequency-independent decrease of intensity across the interface, without significant changes in the line shape.
The perpendicular profile $\tilde{\varepsilon}^{\mathrm{IR}\prime\prime}_{zz}(z)$ (Fig. 5 (d) & (f) & (i)) is more complex. The OH-stretch band, which is in bulk centered around $3500\text{\,}\mathrm{c}\mathrm{m}^{-1}$, exhibits a blue shift when going from bulk to vapor until only the free OH-stretch contribution at $3700\text{\,}\mathrm{c}\mathrm{m}^{-1}$ survives. The libration band [29, 114], centered around $500\text{\,}\mathrm{c}\mathrm{m}^{-1}$, is more pronounced in $\tilde{\varepsilon}^{\mathrm{IR}\prime\prime}_{zz}(z)$ than in $\tilde{\varepsilon}^{\mathrm{IR}\prime\prime}_{xx}(z)$, as can be seen in Figures 5 (e) & (f). This can be rationalized by the higher concentration of planar-oriented molecules near the interface (Fig. 3 (e)), which are more susceptible to $z$-polarized fields.
We observe that isotropically averaged interfacial and bulk spectra are indistinguishable at $z-z_{\mathrm{GDS}}=$-5.5\text{\,}\mathrm{\SIUnitSymbolAngstrom}$\$ in Figure 5 (g) & (j). Hence, the second-order response in Figures 2 & 4, as well as the linear response in Figure 5, and the orientational anisotropy in Figure 3, can be considered bulk-like at this depth. The comparison with experiments [16] (black) in Figure 5 (g) & (j) shows that the differences between simulation and experiments are rather small around the bending peak and in the OH-stretch region at high frequencies ($3500\text{\,}\mathrm{c}\mathrm{m}^{-1}$\$) where the SFG stretch signal is most intense, whereas the experimental IR intensity is significantly underestimated at lower frequencies.
@@ III Conclusion
Accounting for higher-order multipole contributions is essential for interpreting SFG spectra and leads to quantitative agreement between simulations and experimental data. In the bending band, the electric dipole contribution is almost negligible due to the small frequency shift between mutually canceling positive and negative SFG signals from water at different depths, and has an amplitude comparable to the usually neglected magnetic dipole contribution. As a result, the total SFG signal is dominated by the electric quadrupole contribution. The shoulder appearing at $3600\text{\,}\mathrm{c}\mathrm{m}^{-1}$ in the stretch band has been attributed to orientational anisotropy [66, 121], or to a combination mode involving intermolecular coupling [125], as summarized in the review by Tang et al. [128]. Our results demonstrate that the shoulder primarily originates from electric quadrupole contributions, which do not reflect interfacial structure but rather depend on bulk properties for vanishing wave vector mismatch $\Delta k_{z}=0$. Only after subtracting higher-order multipoles can the anisotropic orientation of interfacial molecules be inferred from the SFG signal, by which pronounced biaxial water ordering at the air–water interface is revealed. This biaxial structure consists of essentially three layers of orientational species, namely an inwards-oriented, a planar-oriented, and an outwards-oriented species. This arrangement is surprisingly similar to structural motifs observed at the air-ice interface [94].
In conclusion, since SFG spectra are significantly influenced by non-interface specific multipole effects, it is crucial to account for and subtract these contributions in order to be able to interpret experimental spectra in terms of interfacial orientations. This can be achieved by subtracting a reference spectrum with the same bulk medium [41], or by using theoretical predictions, as those provided in this work.
@@ IV Methods
IV.1 Technical Details
All simulations are done in the NVT ensemble at room temperature $T=$298\text{\,}\mathrm{K}$\$ with the MB-Pol force field [9, 10, 79]. For starting configurations, 94 equally spaced initial configurations are extracted from a $18.8\text{\,}\mathrm{n}\mathrm{s}$ long simulation using the SPC/E force field [14] and the GROMACS molecular dynamics software [1]. The SPC/E simulations are done with a time step of $2\text{\,}\mathrm{f}\mathrm{s}$ and a velocity rescaling thermostat [26], with a relaxation time of $1\text{\,}\mathrm{p}\mathrm{s}$. Each of the initial configurations is equilibrated with the MB-Pol force field for $20\text{\,}\mathrm{p}\mathrm{s}$, and afterward, production runs with an average runtime of $0.86\text{\,}\mathrm{n}\mathrm{s}$ are executed using the LAMMPS software [130]. Again, a velocity-rescaling thermostat is employed. However, for MB-Pol simulations, the relaxation time is set to $5\text{\,}\mathrm{p}\mathrm{s}$, and the time step is $0.2\text{\,}\mathrm{f}\mathrm{s}$. The box size is $2\text{\,}\mathrm{n}\mathrm{m}$ in the x and y dimensions and $6\text{\,}\mathrm{n}\mathrm{m}$ in the z dimension. This size is chosen to ensure that the box contains more than twice as much air as water, since the water slab spans approximately $2.6\text{\,}\mathrm{n}\mathrm{m}$. The box is filled with $352$ water molecules and the two interfaces lie in the $xy$ plane.
The total dipole moment of the simulation box is computed with the modified TTMF-4 model, which describes the electrostatics of the MB-Pol potential [25, 9]. The same model is used to predict the linear absorption profile, where we compute polarization density trajectories from monopole and dipole density trajectories. Molecular polarizabilities are computed with single-molecule quantum chemistry calculations using the software Gaussian 16 [43]. The theoretical level is CCSD(T)/aug-cc-pVTZ and B3LYP/aug-cc-pVTZ in the parameterization of $\alpha^{n,\mathrm{DD}}_{ij}(t)$ and $\alpha^{n,\mathrm{QD}}_{ijk}(t)$, appearing in Equations (22) and (23), respectively. Electrostatic interactions between the induced multipoles are computed as described in SI Sections V B and VIII, involving the Ewald summation algorithm included in OpenMM [35].